OpenStudy (anonymous):

what does d=1/2at2+ViT mean?

OpenStudy (anonymous):

\[distance =\frac{ acceleration \times time ^{2} }{ 2 } +initialvelocity \times time\]

OpenStudy (dls):

\[\LARGE D=ut+\frac{1}{2}at^{2}\] Or \[\LARGE D=V_i t+\frac{1}{2}at^2\] Vi=initial velocity a=acceleration t=time It is one of the Newton's Equation's for uniform motion.